# GaussianFlowOcc: Sparse and Weakly Supervised Occupancy Estimation using Gaussian Splatting and Temporal Flow

### v2

@@ B.2 Rendered Depth and Semantics
In Fig. B.2, we present examples of rendered depth and semantics obtained by applying Gaussian Splatting to the 3D Gaussians estimated from our trained model. For comparison, we also show the pseudo labels used during training, which are generated by GroundedSAM [44] for semantics and Metric3D [63] for depth. The rendered outputs demonstrate a high level of detail, with most objects accurately segmented. Notably, even thin structures such as trees and pedestrians are well-preserved, showcasing the model’s ability to preserve fine geometric details; something voxel-based approaches often struggle with due to grid resolution constraints.
@@
In Fig. C.3, we further illustrate the effect of our Temporal Gaussian Splatting by comparing predicted semantic segmentation maps across consecutive frames: The first row shows the rendered semantics for the current input frame. The second row depicts the segmentation map generated by rendering into the next frame and using the 3D flow estimated by the Temporal Module to compensate the object motion. The third row shows the same next-frame projection but without dynamic object handling — simulating the approach taken by previous works. The last row provides the pseudo-semantic labels of the next frame for comparison. The comparison clearly reveals the importance of dynamic object modeling. Without applying the 3D flow (Row 3), the rendered semantics fail to align with the dynamic objects in the next frame, as these objects have already moved. This misalignment would introduce incorrect supervisory signals if used to compute the loss. In contrast, when applying the Temporal Module’s flow estimation (Row 2), the rendered semantics better match the pseudo labels of the next frame. This demonstrates that our Temporal Module has successfully learned to approximate object motion, improving the alignment of dynamic objects across frames and ensuring more accurate supervision during training.
Figure B.2: Qualitative results on the Occ3D-nuScenes validation set. Each column shows an Input Image, Rendered Depth and Rendered Semantics generated by using GS to render predicted Gaussians into input cameras (as is done during training). We additionally show the pseudo labels used during training for the relevant sample.
@@ Appendix C Voxelization
where $\Sigma_{i}$ represents the covariance matrix of each Gaussian, derived from its rotation quaternion and scale parameters. To ensure efficiency in handling large numbers of Gaussians and voxels, we limit the influence of each Gaussian to a fixed local neighborhood. This is justified by the fact that the contribution of each Gaussian typically decays rapidly with distance, making it computationally unnecessary to evaluate its impact on distant voxels.
Figure C.3: Qualitative results on the Occ3D-nuScenes validation set. Each column shows an input image and the rendered semantics generated by using GS to render predicted Gaussians into the input camera, as well as rendered semantics when rendering the predictions into the next frame, one with using our Temporal Module to compensate motion and one without. The last row displays the pseudo labels of the next frame to show how correct motion compensation should look like.

### v3

@@ B.2 Rendered Depth and Semantics
In LABEL:fig:qual_render, we present examples of rendered depth and semantics obtained by applying Gaussian Splatting to the 3D Gaussians estimated from our trained model. For comparison, we also show the pseudo labels used during training, which are generated by GroundedSAM [45] for semantics and Metric3D [66] for depth. The rendered outputs demonstrate a high level of detail, with most objects accurately segmented. Notably, even thin structures such as trees and pedestrians are well-preserved, showcasing the model’s ability to preserve fine geometric details; something voxel-based approaches often struggle with due to grid resolution constraints.
@@
In LABEL:fig:qual_render_temporal, we further illustrate the effect of our Temporal Gaussian Splatting by comparing predicted semantic segmentation maps across consecutive frames: The first row shows the rendered semantics for the current input frame. The second row depicts the segmentation map generated by rendering into the next frame and using the 3D flow estimated by the Temporal Module to compensate the object motion. The third row shows the same next-frame projection but without dynamic object handling — simulating the approach taken by previous works. The last row provides the pseudo-semantic labels of the next frame for comparison. The comparison clearly reveals the importance of dynamic object modeling. Without applying the 3D flow (Row 3), the rendered semantics fail to align with the dynamic objects in the next frame, as these objects have already moved. This misalignment would introduce incorrect supervisory signals if used to compute the loss. In contrast, when applying the Temporal Module’s flow estimation (Row 2), the rendered semantics better match the pseudo labels of the next frame. This demonstrates that our Temporal Module has successfully learned to approximate object motion, improving the alignment of dynamic objects across frames and ensuring more accurate supervision during training.
B.3 Long-term Temporal Flow
We further examine the behavior of the Temporal Module by visualizing the estimated flow over a longer temporal horizon, spanning three frames into the past and future. Figure B.2 displays rendered semantic maps that illustrate the qualitative effectiveness of the estimated 3D flow. As is always the case, the model predicts the 3D Gaussians for the current frame $t$ given the images of the current frame (along with the previous frame’s Gaussians for temporal fusion). The Temporal Module then estimates the motion of objects across each of the surrounding frames, and the Gaussian means are updated according to the predicted flow. The figure compares rendered results using Gaussians moved by the estimated flow with those using static, unmoved Gaussians. The visualizations demonstrate that the model learns reasonable object motion that better aligns with the temporal position of the camera frame, despite not being explicitly trained with motion or scene flow supervision. The Temporal Module is optimized solely through consistency constraints between rendered temporal frames and the Gaussians adjusted by the estimated flow. This temporal consistency reduces artifacts during Temporal Gaussian Splatting, which is critical for stable training. We note that learning scene flow in this weakly supervised manner remains challenging, and errors are still present in practice. Nevertheless, as shown in Sec. 4.4.3, leveraging the estimated flow to improve temporal consistency leads to a notable increase in occupancy estimation performance.
Figure B.2: Qualitative results on the Occ3D-nuScenes validation set. Each row depicts the rendered estimated Gaussians when using the estimated temporal flow to move the Gaussians versus not using the flow, for a time horizon of 1.5s (3 frames) into the past and the future, respectively. The Gaussians are predicted for the current frame, and then rendered into adjacent frames. It is visible that the model learns to compensate object motion over multiple frames.
B.4 Comparison to GaussianOcc
To further demonstrate the benefits of our approach, we compare the predicted occupancy between GaussianFlowOcc and GaussianOcc [15] in Fig. B.3. In contrast to GaussianOcc, our model consistently detects and models thin and small objects like trees, traffic signs and poles (scene-0916 & scene-0904). In addition, GaussianFlowOcc suffers a lot less from object bleeding (scene-0557 & scene-0775). It is visible that our method can much better estimate the 3D shape of objects (such as vehicles) avoiding the characteristic depth stretching that often affects 2D-supervised methods [41, 15, 4, 20, 70]. We attribute this improvement to our use of long-term temporal supervision, which enforces geometric consistency over time. This strategy is only viable with an explicit dynamic scene model—highlighting the importance of our integrated motion estimation framework.
Figure B.3: Comparing our method to GaussianOcc. We compare the occupancy estimation results on four samples of different validation scenes of the nuScenes dataset between our method and GaussianOcc [15]. It is visible that our method can better represent thin and small objects, suffers less from object bleeding and can better estimate the complete 3D shape of scene objects.
@@ Appendix C Voxelization
where $\Sigma_{i}$ represents the covariance matrix of each Gaussian, derived from its rotation quaternion and scale parameters. To ensure efficiency in handling large numbers of Gaussians and voxels, we limit the influence of each Gaussian to a fixed local neighborhood. This is justified by the fact that the contribution of each Gaussian typically decays rapidly with distance, making it computationally unnecessary to evaluate its impact on distant voxels. As explained in Sec. A.3, this voxelization operation is fully differentiable, which enables training of our model using 3D voxel labels when available.
Appendix D Source Code
Source code to reproduce the results presented in the paper is available at https://github.com/boschresearch/GaussianFlowOcc. Please follow the instructions in the README.md file to install the repository and run the code. We provide instructions to train and evaluate our models on the Occ3D-nuScenes dataset.